OpenStudy (anonymous):

Please help. I'm desperate. For the graphed function f(x) = (2)^x + 2 + 1, calculate the average rate of change from x = 0 to x = 2. options are: -3/2 -2 3/2 4

OpenStudy (mathmale):

The formula for "average rate of change" is much like the formula for slope (m). I'll type it out for you:\[ave~rate~of~change=\frac{ f(b)-f(a) }{ b-a }\]Would you please share whatever you know about this formula. Where do a and b come from? What does f(x) represent?

OpenStudy (anonymous):

I really know nothing. I slacked off in the beginning of the year. So I really just need the answer. I'm desperate. I'm freaking out cause i might fail the class. Please help

OpenStudy (anonymous):

heres the picture that came with it

OpenStudy (anonymous):

@mathgeek1 @mathmale please help

OpenStudy (mathmale):

I'd be glad to help, but I do expect you to do some appropriate preparation before posting these questions. Sorry you slacked off earlier this term; what you are facing now are direct consequences of that slacking off. share whatever you do know about each problem. then I'd be happy to step in with suggestions and guidance.

OpenStudy (anonymous):

Im being serious. Im looking at these problems and drawing a complete blank. I know absaloutly nothing in this worksheet

OpenStudy (anonymous):

i havea low grade in the class already. i need to get a good grade on this. please help

OpenStudy (mathmale):

All I'm asking is that you demonstrate SOME preparation, some basic groundwork; if you haven't done any prep at all, then it's awfully late to be asking for help. Again, please ask some basic questions so I can respond meaningfully. I gave you a formula for the average rate of change of a function. Have you seen that before? Basically, to find the average rate of change of a function from x=a to x=b, we evaluate the function at x=a and at x=b, subtract f(a) from f(b), and then divide that result by the change in x, which is just b-a. Here your a= 0 and your b = 2.

OpenStudy (mathmale):

Appeals to your being desperate won't help much here. Sorry to sound tough. I just want you involved in the problem solving process.

OpenStudy (anonymous):

i've researched everything already trying to understand, but i just cant. could you please just help me answer this. i know i shouldve payed attention but in order for me to understand this stuff i'd have to retake my entire high school career. please just help

OpenStudy (mathmale):

You are given the function f(x) = (2)^x + 2 + 1. You are told that x changes from x=a=0 to x=b=2. Evaluate f(x) at x=0. Evaluate f(x) at x=2.

OpenStudy (anonymous):

i dont know how!

OpenStudy (mathmale):

Hint: In f(x) = (2)^x + 2 + 1, throw out every occurrence of x and replace it with 0. f(0)=?

OpenStudy (anonymous):

is the answer -2?

OpenStudy (mathmale):

Were f(x) = (2)^x + 2 + 1 correct, you'd get f(0)=2^0+2+1, or 1+2+1. What's that equal to?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

thats the answer?

OpenStudy (mathmale):

do you know how to take and share screen shots?

OpenStudy (anonymous):

yes. but is 4 the answer?

OpenStudy (mathmale):

Before I answer that, please take a screen shot of the original problem and upload it herer. I suspect that the last two terms in f(x) = (2)^x + 2 + 1 are not correct. Note that 2+1=3, and that's suspicious looking (too simple).

OpenStudy (anonymous):

its to the left of the screen

OpenStudy (mathmale):

thank you. Sorry to bring you bad news, but your function is not f(x) = (2)^x + 2 + 1. Instead, it's\[f(x)=2^{x+2}\] which is quite different from f(x) = (2)^x + 2 + 1. Next time, write f(x) this way: f(x) = (2)^(x + 2) + 1. Please evaluate f(x) = (2)^(x + 2) + 1. at x=0. Note that (0+2)=2

OpenStudy (anonymous):

so what would be the answser

OpenStudy (mathmale):

Maddie, I'm asking you that question. I know how to find the answer; my job is to help YOU to find it for yourself.

OpenStudy (anonymous):

its 4?

OpenStudy (mathmale):

f(x) = (2)^(x + 2) + 1, evaluated at f(0), is f(0)=(2)^2 + 1 = 4+1 = ????

OpenStudy (anonymous):

i mean 5

OpenStudy (mathmale):

f(0)=(2)^2 + 1 = 4+1 = ????

OpenStudy (anonymous):

i know but 5 isnt an option

OpenStudy (mathmale):

that's better. f(0)=5. Now find f(2). Wherever you see x in the function f(x)=(2)^(x+2) + 1 , substitute 2. f(2)=(2)^(2+2) + 1 = ???

OpenStudy (anonymous):

32

OpenStudy (mathmale):

But 2^4=16 and 16+1=??

OpenStudy (mathmale):

f(2)=(2)^(2+2) + 1 => (2)^4 = 1 = 16 + 1 = ??

OpenStudy (anonymous):

17

OpenStudy (anonymous):

is it 17? whats the next step

OpenStudy (mathmale):

Now you have f(0), f(2), 0 and 2; Evaluate \[\frac{ f(2)-f(0) }{ 2-0 }\]

OpenStudy (anonymous):

f(2)/2

OpenStudy (mathmale):

f(2) = 17 f(0) = 5 so f(2) - f(0) = 17 = 5 = 12. a = 0 b = 2 so b - a = ??

OpenStudy (anonymous):

2

OpenStudy (anonymous):

so the answer is -2?

OpenStudy (mathmale):

f(2) - f(0) is 17 = 5 = 12. Next, 2-0 is 2. Next, 12/2 is 6. the average rate of change of your function 2^(x+2) + 1 over the interval ]0,2} is thus 6. THAT is the answer you wanted.

OpenStudy (mathmale):

Review: Average Rate of Change of f(x) over an interval [a,b] is\[A.R.C.=\frac{ f(b)-f(a) }{ b-a }\]

OpenStudy (mathmale):

and here your A. R. C. is\[A.R.C.=\frac{ f(2)-f(0) }{ 2-0 }=\frac{ 17-5 }{ 2 }=\frac{ 12 }{ 2 }=6\]